# On the Effectiveness of Code Representation in Deep Learning-Based Automated Patch Correctness Assessment

### 08 Mar 2026 v1

@@ 1. Introduction
During the maintenance and evolution of modern software systems, software bugs are inevitable and may result in fatal consequences (Wang et al., 2017; Zhang et al., 2022a). It is increasingly time-consuming and labor-intensive for developers to manually fix detected bugs due to the rising number and complexity of software systems (Hossain, 2018). For example, prior work (Britton et al., 2013) demonstrates that developers usually spend 50% of their time on software debugging and fixing. To relieve manual debugging efforts, Automated Program Repair (APR) has been proposed to generate correct patches automatically without human intervention, and has received a great deal of attention from both academia and industry (Gazzola et al., 2019; Zhang et al., 2023a). In the past decades, a mass of APR approaches have been proposed and have made substantial progress on the number of correctly-fixed bugs, including heuristic-based, constraint-based, pattern-based, and learning-based ones (Monperrus, 2018; Le Goues et al., 2012; Liu et al., 2019b; Li et al., 2020; Yuan et al., 2022).
However, the APR community has fundamentally been challenged by the patch overfitting issue due to insufficient program specifications (Smith et al., 2015; Le et al., 2019; Liang et al., 2021; Motwani et al., 2022; Motwani and Brun, 2023). In particular, existing APR approaches typically adopted the generate-and-validate paradigm, i.e., generating candidate patches and leveraging the developer-written test cases to identify correct patches (Zhang et al., 2023a). In such cases, candidate patches passing the available test cases are returned, but they may not be generalized to other potential test cases, leading to overfitting patches. Previous studies (Tao et al., 2014; Zhang et al., 2022b) demonstrate that it may take considerable time and effort for developers to filter out the overfitting patches, hindering the practical usability of off-the-shelf APR tools in real-world debugging scenarios.
@@
This Paper. Despite an emerging research area, the APCA community has seen different well-designed code representation approaches to encode patched code snippets, e.g., heuristic-based (Ye et al., 2022), sequence-based (Tian et al., 2020), and tree-based features (Lin et al., 2022). Code representation is crucial and fundamental in learning the program semantics and reasoning about patch correctness, and has already been widely investigated in various code understanding tasks (Siow et al., 2022; Sun et al., 2023; Namavar et al., 2022; Utkin et al., 2022). However, the literature does not yet provide a systematic evaluation of how to represent source code for the patch correctness assessment task. There exist several unsolved challenges about code representations in APCA. (1) What type of code representation is optimal in the patch correctness assessment scenario? (2) Despite the superiority shown by heuristic-based, sequence-based, and tree-based approaches, the graph-based approach remains underexplored. (3) Can such different representations be integrated to further enhance the prediction effectiveness? In this paper, to fill this gap, we investigate the feasibility of leveraging different code representation approaches to predict patch correctness and facilitate subsequent APCA studies. Based on the above challenges, we address the following three research questions (RQs): RQ1 analyzes the impact of different representations; RQ2 answers the comparison performance with state-of-the-art approaches; and RQ3 explores the potential of fusion-based approaches.
@@
Open Science. To facilitate future research in the APCA community, we release the replication artifacts, including the studied dataset, source code, model training and evaluation scripts, and related models in our experiment (Zhang and Hu, 2024).
@@ 2.1.1. Repair Workflow
Given a buggy program and some available test cases that make the program fail, the primary objective of APR is to generate patches automatically that pass all available test suites. Existing APR techniques typically follow the generate-and-validate workflow, as illustrated in Fig. 1. In particular, for a detected bug, (1) the fault localization phase identifies suspicious code elements and returns the code snippets that need to be patched; (2) the patch generation phase generates candidate patches by performing program transformation while keeping the program semantically correct. (3) the patch validation phase filters out incorrect patches by executing available test suites against all generated patches. In the literature, a mass of APR techniques have been proposed to generate patches from different aspects. APR techniques can be categorized into four classes, including heuristic-based (Le Goues et al., 2012; Martinez and Monperrus, 2016; Yuan and Banzhaf, 2018), constraint-based (Martinez and Monperrus, 2018; Durieux and Monperrus, 2016; Mechtaev et al., 2016), pattern-based (Koyuncu et al., 2020; Liu et al., 2019a, b; Le et al., 2016) and learning-based (Tufano et al., 2019; Zhu et al., 2021; Ye and Monperrus, 2024) ones. We detail them in Section 6.1.
@@ 2.1.2. Patch Overfitting Issue
During patch generation, a candidate patch passing the original test suite is called a plausible patch. A plausible patch, which is also semantically equivalent to the developer patch, denotes a correct patch. However, existing test suites are inherently incomplete to cover the entire behavioral domain of a program. Thus, some plausible patches passing available test suites may not be generalized to potential test suites, resulting in a critical challenge in APR literature, i.e., the patch overfitting issue. For example, Qi et al. (Qi et al., 2015) find that a majority of plausible patches generated by APR approaches are overfitting the available test suite while actually being incorrect. When provided to developers, such overfitting patches require extensive manual inspection efforts and lead to negative debugging performance (Tao et al., 2014; Zhang et al., 2022b), limiting the usability of existing APR approaches in practice.
@@ 2.2. Code Representation
With the success of DL in the SE community, code representation has been foundational for learning the semantics of source code to facilitate code-related task automation (Watson et al., 2022; Yang et al., 2022; Zhang et al., 2023b). According to previous SE studies (Zhang et al., 2023a; Siow et al., 2022; Sun et al., 2023), code representation can be categorized into four types, including heuristic-based, sequence-based, tree-based, and graph-based representations. Heuristic-based representation refers to representing source code as a set of hand-crafted features, such as frequency-inverse document frequency (TF-IDF). Such features are usually pre-designed by domain experts to identify important aspects of source code based on specific tasks, and utilized in various tasks such as patch correctness assessment (Ye et al., 2022) and fault localization (Wong et al., 2016). Sequence-based representation treats source code as a sequence of tokens, which is similar to natural language. A mass of studies employ sequence-based representation to address code-related tasks, such as program repair (Yuan et al., 2022; Zhang et al., 2024b) and patch correctness assessment (Tian et al., 2020, 2023; Zhang et al., 2024a), demonstrating that the sequence-based representation is capable of learning semantic and syntactic information of source code. Tree-based representation attempts further to capture the structure information of source code with AST. Such representation is widely-used in previous studies, such as program repair (Li et al., 2020, 2022) and patch correctness assessment (Lin et al., 2022). Graph-based representation further considers some semantic information that is not well-modeled in tree-based representation, such as control flow and data flow. Researchers have proposed various approaches with graph-based representation for source code, such as program repair (Zhu et al., 2021, 2023) and vulnerability detection(Li et al., 2021; Zhou et al., 2019). In this work, we consider all four types of code representations to train classification models for patch correctness prediction.
@@ 3. Study Design
Figure 2. Overview of our empirical study
@@ 3.1. Overview
Overall, our work aims to investigate the effectiveness of various code representations in reasoning about patch correctness. Fig. 2 presents a typical prediction pipeline of automated patch correctness assessment when DL models are incorporated. The workflow comprises three primary phases, illustrated as follows. First, given a plausible patch generated by off-the-shelf APR tools and passing available functional test cases, we perform the input preprocessing phase to extract its buggy code snippet and patched code snippet. Second, we represent the buggy and patched code snippet from different perspectives to capture their features, including heuristic-based, sequence-based, tree-based and graph-based strategies. Third, the extracted code features are used to train DL models to predict patch correctness.
@@ 3.3.1. Heuristic-based Representation
We first consider TF-IDF to vectorize code snippets, which has been widely adopted in previous studies (Siow et al., 2022). In addition, Following ODS (Ye et al., 2022), we adopt all 202 hand-crafted features, which are categorized into three types (i.e., code, context, and pattern) as per the original design of ODS.
@@ 4. Evaluation and Results
4.1. Answer to RQ1: Comparison with Different Code Representations
@@
Results of Tree-based Representation. Table 4 presents the comparison results of tree-based representation across five metrics and two models. We can find that TreeLSTM demonstrates superior performance over TBCNN across all the evaluated metrics. Specifically, TreeLSTM achieves 82.94% for accuracy, 80.01% for precision, 86.73% for recall, 83.18% for F1, and 83.03% for AUC. The improvement against TBCMM reaches 1.49%, 0.97%, 2.49%, 1.64%, and 1.52%, indicating the ability of TreeLSTM to identify more overfitting patches with a lower rate of false positives.
@@
Results of Graph-based Representation. Table 5 presents the comparison results of graph-based representation across five graph-based representations, three models, and five metrics. When comparing different graphs, we find CPG achieves optimal performance, with accuracy, precision, recall, F1, and AUC being 82.69%, 79.27%, 87.45%, 83.12%, 82.80%, and 83.07% on average across three classifiers. The overall improvement of CPG against CFG, DDG, PDG, and CDG across five metrics reaches 3.65%, 2.90%, 3.84%, and 7.53%, respectively. Such improvement of CPG is reasonable, as CPG is a combination of several graphs, thereby containing more information, such as data dependency and control dependency. The second-best performance is exhibited by CFG, DDG, and PDG, which attain comparable performance, e.g., 80.15%, 80.72%, and 80.00% for five metrics on average. Besides, CDG performs the worst among the five graphs with an accuracy of only 76.56% on average, which is 2.90%, 3.47%, 2.75%, and 5.82% less than CFG, DDG, PDG, and CPG. When comparing the three models, we find that GGNN yields consistently optimal performance and GCN achieves the second-best performance across all graph types and metrics. The improvement of GGNN against GCN reaches 10.43%, 11.04%, 9.46%, 12.78%, and 1.74%, respectively. The possible reason lies in that GGNN utilizes GRU to update node representations in the graph, which allows for the incorporation of information from a node’s neighbors. Similarly, GCN outperforms GAT by 2.45%, 1.97%, 1.39%, 7.92% and 1.30% on the five metrics.
Comparison of Four Types of Representation. From Tables 2 to 5, we first find that sequence-based approaches have a better performance compared with the heuristic-based approaches. with an improvement of 2.57% for accuracy and 4.88% for recall. Second, the difference between the sequence-based and tree-based approaches is not very obvious, e.g., only 0.41% improvement. Third, we find that the graph-based approach achieves the best performance compared with other representations, with an average of 79.99% for accuracy, 78.36% for precision, 82.54% for recall, 80.31% for F1, and 80.00% for AUC across all graphs and classifiers. For example, GGNN with CGP outperforms XFBoost with TF-IDF, transformer, and TreeLSTM by 4.19%, 1.29%, and 0.88% on average, indicating that the semantic and syntactic information on graphs are beneficial in learning the semantic of the buggy program.
@@
4.2. Answer to RQ2: Comparison with State-of-the-art Approaches
@@
4.3. Answer to RQ3: Effectiveness of Fusion-based APCA Approach
Experimental Design. As mentioned in Section 4.1, different representations are able to complement each other in identifying overfitting patches. Different representations capture various aspects of source code, such as semantic and structural information. Thus, in this section, we investigate the potential of combining these features to improve the prediction performance. We consider two strategies to combine different representations according to the stage of fusion, i.e., intermediate fusion and backend fusion. The first one attempts to concatenate feature vectors for training prediction models, and the second one performs ensemble learning on already trained prediction models.
@@ 4.3. Answer to RQ3: Effectiveness of Fusion-based APCA Approach
First, we find that, fusion is able to further boost heuristic-based and sequence-based representations. For example, for the intermediate fusion, by combining sequence-based, tree-based, and graph-based representations, heuristic-based representation is improved by 11.97%, 10.81%, and 11.80% on five metrics, respectively. Similarly, for the backend fusion, the improvement of heuristic-sequence, heuristic-tree, and heuristic-graph representations against the single heuristic-based representation achieves 13.58%, 8.77%, and 10.28% on average. However, for tree-based and graph-based representations, fusion fails to always bring better results. For example, the accuracy of the tree-based representation is improved by 1.62% with the integration of the heuristic-based representation, while being decreased by 0.01% with the integration of the heuristic-based representation. The possible reason may be that tree-based and graph-based representations have already achieved impressive performance before fusion, which is difficult to be further improved compared with two relatively poor-performing representations, i.e., heuristic-based and sequence-based representations.
Second, we find that the combination of more than two representations does not yield further enhancements. For example, heuristic-sequence-tree is able to achieve a prediction accuracy of 82.10%, outperforming sequence-tree by 0.48%, while underperforming heuristic-sequence and heuristic-tree by 2.49% and 1.13%. The possible reason is that the fusion methods (i.e., the vector concatenation in intermediate fusion and the average prediction probability in backend fusion) employed are naive to extract the deep relationships among the various representations. Thus, future researchers should further explore advanced techniques that can fusion various patch representations for better prediction performance. Third, when comparing the two fusion strategies, we find that, the intermediate fusion outperforms the backend fusion by 1.22% of accuracy, 0.75% of precision, 1.70% of recall, 1.23% of F1, 1.20% of AUC on the heuristic-sequence representation. However, for the other two representation fusion scenarios, the backend fusion achieves better or slightly better results than the intermediate fusion, such as an average improvement of 1.55% 1.15%, 1.28%, 0.69%, and 0.78% on the heuristic-tree, heuristic-graph, sequence-tree, sequence-graph, tree-graph representations. Besides, for the fusion of more than two features, the two fusion methods achieve comparable results, such as 0.36%, 0.39%, 0.20%, 0.17%, 0.71% improvement or disimprovement on five representations, as listed in the last five rows in Table 7.
Answer to RQ3: Our comparison results demonstrate that, the fusion of two representations can improve the prediction performance, while incorporating more representations tends to have a detrimental effect due to the simplistic nature of the utilized fusion models.

### 05 Apr 2026 v2

@@ 1. Introduction
During the maintenance and evolution of modern software systems, software bugs are inevitable and may result in fatal consequences (Gazzola et al., 2019). It is increasingly time-consuming and labor-intensive for developers to manually fix detected bugs due to the rising number and complexity of software systems (Hossain, 2018). For example, prior work (Britton et al., 2013) demonstrates that developers usually spend 50% of their time on software debugging and fixing. To relieve manual debugging efforts, Automated Program Repair (APR) has been proposed to generate correct patches automatically without human intervention, and has received a great deal of attention from both academia and industry (Zhang et al., 2023a, ). In the past decades, a mass of APR approaches have been proposed and have made substantial progress on the number of correctly-fixed bugs, including heuristic-based, constraint-based, pattern-based, and learning-based ones (Monperrus, 2018; Le Goues et al., 2012; Liu et al., 2019b; Li et al., 2020; Xia et al., 2025).
However, the APR community has fundamentally been challenged by the patch overfitting issue due to insufficient program specifications (Smith et al., 2015; Le et al., 2019; Liang et al., 2021; Motwani et al., 2022; Motwani and Brun, 2023). In particular, existing APR approaches typically adopted the generate-and-validate paradigm, i.e., generating candidate patches and leveraging the developer-written test cases to identify correct patches (Zhang et al., 2023a). In such cases, candidate patches passing the available test cases are returned, but they may not be generalized to other potential test cases, leading to overfitting patches. Previous studies (Wang et al., 2025; Tao et al., 2014) demonstrate that it may take considerable time and effort for developers to filter out the overfitting patches, hindering the practical usability of off-the-shelf APR tools in real-world debugging scenarios.
@@
This Paper. Despite an emerging research area, the APCA community has seen different well-designed code representation approaches to encode patched code snippets, e.g., heuristic-based (Ye et al., 2022), sequence-based (Tian et al., 2020), and tree-based features (Lin et al., 2022). Code representation is crucial and fundamental in learning the program semantics and reasoning about patch correctness, and has already been widely investigated in various code understanding tasks (Siow et al., 2022; Sun et al., 2026; Namavar et al., 2022; Utkin et al., 2022). However, the literature does not yet provide a systematic evaluation of how to represent source code for the patch correctness assessment task. There exist several unsolved challenges about code representations in APCA. (1) What type of code representation is optimal in the patch correctness assessment scenario? (2) Despite the superiority shown by heuristic-based, sequence-based, and tree-based approaches, the graph-based approach remains underexplored. (3) Can such different representations be integrated to further enhance the prediction effectiveness? In this paper, to fill this gap, we investigate the feasibility of leveraging different code representation approaches to predict patch correctness and facilitate subsequent APCA studies. Based on the above challenges, we address the following three research questions (RQs): RQ1 analyzes the impact of different representations; RQ2 answers the comparison performance with state-of-the-art approaches; and RQ3 explores the potential of fusion-based approaches.
@@
Open Science. To facilitate future research in the APCA community, we release the replication artifacts, including the studied dataset, source code, model training and evaluation scripts, and related models in our experiment (Authors, 2026).
@@ 2.1.1. Repair Workflow
Given a buggy program and some available test cases that make the program fail, the primary objective of APR is to generate patches automatically that pass all available test suites. Existing APR techniques typically follow the generate-and-validate workflow, as illustrated in Fig. 1. In particular, for a detected bug, (1) the fault localization phase identifies suspicious code elements and returns the code snippets that need to be patched; (2) the patch generation phase generates candidate patches by performing program transformation while keeping the program semantically correct; (3) the patch validation phase filters out incorrect patches by executing available test suites against all generated patches. In the literature, a mass of APR techniques have been proposed to generate patches from different aspects. APR techniques can be categorized into four classes, including heuristic-based (Le Goues et al., 2012; Martinez and Monperrus, 2016; Yuan and Banzhaf, 2018), constraint-based (Martinez and Monperrus, 2018; Durieux and Monperrus, 2016; Mechtaev et al., 2016), pattern-based (Koyuncu et al., 2020; Liu et al., 2019a, b; Le et al., 2016) and learning-based (Tufano et al., 2019; Zhu et al., 2021; Ye and Monperrus, 2024) ones. We detail them in Section 6.1.
@@ 2.1.2. Patch Overfitting Issue
During patch generation, a candidate patch passing the original test suite is called a plausible patch. A plausible patch, which is also semantically equivalent to the developer patch, denotes a correct patch. However, existing test suites are inherently incomplete to cover the entire behavioral domain of a program. Thus, some plausible patches passing available test suites may not be generalized to potential test suites, resulting in a critical challenge in APR literature, i.e., the patch overfitting issue. For example, Qi et al. (Qi et al., 2015) find that a majority of plausible patches generated by APR approaches are overfitting the available test suite while actually being incorrect. When provided to developers, such overfitting patches require extensive manual inspection efforts and lead to negative debugging performance (Tao et al., 2014; Zhang et al., 2022), limiting the usability of existing APR approaches in practice.
@@ 2.2. Code Representation
With the success of DL in the SE community, code representation has been foundational for learning the semantics of source code to facilitate code-related task automation (Watson et al., 2022; Yang et al., 2022; Zhang et al., 2026a). According to previous SE studies (Zhang et al., 2023a; Siow et al., 2022; Sun et al., 2026), code representation can be categorized into four types, including heuristic-based, sequence-based, tree-based, and graph-based representations. heuristic-based representation refers to representing source code as a set of hand-crafted features, such as frequency-inverse document frequency (TF-IDF). Such features are usually pre-designed by domain experts to identify important aspects of source code based on specific tasks, and utilized in various tasks such as patch correctness assessment (Ye et al., 2022) and fault localization (Wong et al., 2016). Sequence-based representation treats source code as a sequence of tokens, which is similar to natural language. A mass of studies employ sequence-based representation to address code-related tasks, such as program repair (Yuan et al., 2022; Zhang et al., 2024b) and patch correctness assessment (Tian et al., 2020, 2023; Zhang et al., 2024a), demonstrating that the sequence-based representation is capable of learning semantic and syntactic information of source code. Tree-based representation attempts further to capture the structure information of source code with AST. Such representation is widely-used in previous studies, such as program repair (Li et al., 2020, 2022) and patch correctness assessment (Lin et al., 2022). Graph-based representation further considers some semantic information that is not well-modeled in tree-based representation, such as control flow and data flow. Researchers have proposed various approaches with graph-based representation for source code, such as program repair (Zhu et al., 2021, 2023) and vulnerability detection(Li et al., 2021; Zhou et al., 2019). In this work, we consider all four types of code representations to train classification models for patch correctness prediction.
@@ 3. Study Design
Figure 2. Overview of our study
@@ 3.1. Overview
Overall, our work aims to investigate the effectiveness of various code representations in reasoning about patch correctness. Fig. 2 presents a typical prediction pipeline of automated patch correctness assessment when DL models are incorporated. The workflow comprises three primary phases, illustrated as follows. First, given a plausible patch generated by off-the-shelf APR tools and passing available functional test cases, we perform the input preprocessing phase to extract the buggy code snippet and the patched code snippet. Second, we represent the buggy and patched code snippet from different perspectives to capture their features, including heuristic-based, sequence-based, tree-based and graph-based strategies. Third, the extracted code features are used to train DL models to predict patch correctness.
@@ 3.3. Representation Selection
We first consider TF-IDF for vectorizing code snippets, which has been widely adopted in prior studies (Siow et al., 2022). In addition, Following ODS (Ye et al., 2022), we adopt all 202 hand-crafted features, which are categorized into three types (i.e., code, context, and pattern) as per the original design of ODS.
@@ 4. Evaluation and Results
4.1. RQ1: Comparison with Different Code Representations
@@
Results of Tree-based Representation. Table 4 presents the comparison results of tree-based representation across five metrics and two models. We find that TreeLSTM outperforms TBCNN across all evaluated metrics. Specifically, TreeLSTM achieves 82.94% accuracy, 80.01% precision, 86.73% recall, 83.18% F1, and 83.03% AUC. The improvement over TBCMM is 1.49%, 0.97%, 2.49%, 1.64%, and 1.52%, indicating TreeLSTM’s ability to identify more overfitting patches with a lower false-positive rate.
@@
Results of Graph-based Representation. Table 5 presents the comparison results of graph-based representation across five graph-based representations, three models, and five metrics. When comparing different graphs, we find CPG achieves optimal performance, with accuracy, precision, recall, F1, and AUC being 82.69%, 79.27%, 87.45%, 83.12%, 82.80%, and 83.07% on average across three classifiers. The overall improvement of CPG against CFG, DDG, PDG, and CDG across five metrics reaches 3.65%, 2.90%, 3.84%, and 7.53%, respectively. Such improvement of CPG is reasonable, as CPG is a combination of several graphs, thereby containing more information, such as data dependency and control dependency. The second-best performance is exhibited by CFG, DDG, and PDG, which attain comparable performance, e.g., 80.15%, 80.72%, and 80.00% for five metrics on average. Besides, CDG performs the worst among the five graphs with an accuracy of only 76.56% on average, which is 2.90%, 3.47%, 2.75%, and 5.82% less than CFG, DDG, PDG, and CPG. When comparing the three models, we find that GGNN yields consistently optimal performance and GCN achieves the second-best performance across all graph types and metrics. The improvement of GGNN against GCN reaches 10.43%, 11.04%, 9.46%, 12.78%, and 1.74%, respectively. The possible reason lies in the fact that GGNN utilizes GRU to update node representations in the graph, which allows for the incorporation of information from a node’s neighbors. Similarly, GCN outperforms GAT by 2.45%, 1.97%, 1.39%, 7.92% and 1.30% on the five metrics.
Comparison of Four Types of Representation. From Tables 2 to 5, we first find that sequence-based approaches have a better performance compared with the heuristic-based approaches. with an improvement of 2.57% for accuracy and 4.88% for recall. Second, the difference between the sequence-based and tree-based approaches is not very obvious, e.g., only 0.41% improvement. Third, we find that the graph-based approach achieves the best performance among representations, with an average accuracy of 79.99%, a precision of 78.36%, a recall of 82.54%, an F1 of 80.31%, and an AUC of 80.00% across all graphs and classifiers. For example, GGNN with CGP outperforms XFBoost with TF-IDF, transformer, and TreeLSTM by 4.19%, 1.29%, and 0.88% on average, indicating that the semantic and syntactic information on graphs is beneficial in learning the semantics of the buggy program.
@@
4.2. RQ2: Comparison with State-of-the-art Approaches
@@
4.3. RQ3: Effectiveness of Fusion-based APCA Approach
Experimental Design. As mentioned in Section 4.1, different representations are able to complement each other in identifying overfitting patches. Different representations capture various aspects of source code, such as semantic and structural information. Thus, in this section, we investigate the potential of combining these features to improve the prediction performance. We consider two strategies to combine different representations according to the stage of fusion, i.e., intermediate fusion and backend fusion. The first one attempts to concatenate feature vectors to train prediction models, and the second one performs ensemble learning on already trained prediction models.
@@ 4.3. RQ3: Effectiveness of Fusion-based APCA Approach
First, we find that fusion can further boost heuristic-based and sequence-based representations. For example, for the intermediate fusion, by combining sequence-based, tree-based, and graph-based representations, the heuristic-based representation is improved by 11.97%, 10.81%, and 11.80% on five metrics, respectively. Similarly, for the backend fusion, the improvement of heuristic-sequence, heuristic-tree, and heuristic-graph representations against the single heuristic-based representation achieves 13.58%, 8.77%, and 10.28% on average. However, for tree-based and graph-based representations, fusion fails to always bring better results. For example, the accuracy of the tree-based representation is improved by 1.62% with the integration of the heuristic-based representation, while being decreased by 0.01% with the integration of the heuristic-based representation. The possible reason may be that tree-based and graph-based representations have already achieved impressive performance before fusion, which is difficult to be further improved compared with two relatively poor-performing representations, i.e., heuristic-based and sequence-based representations.
Second, we find that combining more than two representations does not yield further improvements. For example, heuristic-sequence-tree is able to achieve a prediction accuracy of 82.10%, outperforming sequence-tree by 0.48%, while underperforming heuristic-sequence and heuristic-tree by 2.49% and 1.13%. The possible reason is that the fusion methods (i.e., the vector concatenation in intermediate fusion and the average prediction probability in backend fusion) employed are naive to extract the deep relationships among the various representations. Thus, future researchers should further explore advanced techniques that can fuse various patch representations for better prediction performance. Third, when comparing the two fusion strategies, we find that the intermediate fusion outperforms the backend fusion by 1.22% of accuracy, 0.75% of precision, 1.70% of recall, 1.23% of F1, and 1.20% of AUC on the heuristic-sequence representation. However, for the other two representation fusion scenarios, the backend fusion achieves better or slightly better results than the intermediate fusion, with average improvements of 1.55%, 1.15%, 1.28%, 0.69%, and 0.78% on the heuristic-tree, heuristic-graph, sequence-tree, sequence-graph, and tree-graph representations. Besides, for the fusion of more than two features, the two fusion methods achieve comparable results, such as 0.36%, 0.39%, 0.20%, 0.17%, 0.71% improvement or disimprovement on five representations, as listed in the last five rows in Table 7.
Answer to RQ3: Our comparison results demonstrate that the fusion of two representations can improve the prediction performance, while incorporating more representations tends to have a detrimental effect due to the simplistic nature of the utilized fusion models.